OpenStudy (anonymous):

What value of m makes the statement true? m + (–9.5) = –7.8

OpenStudy (anonymous):

@RavenTeenTitans

OpenStudy (anonymous):

you take the opposite approach. ((instead of subtract, add.))

OpenStudy (solomonzelman):

add 9.5 to both sides

OpenStudy (anonymous):

um...

OpenStudy (anonymous):

1.7

OpenStudy (anonymous):

its 1.7...I think.....

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

\(\large\color{black}{ m + (–9.5) = –7.8 }\) \(\large\color{black}{ m –9.5 = –7.8 }\) \(\large\color{black}{ m –9.5\color{magenta}{+9.5}= –7.8 \color{magenta}{+9.5}}\) \(\large\color{black}{ m = 1.7 }\)